Q2. Solve xy + y + 9xy = 0 in terms of Bessel functions. Q3. Consider the power series S. (x) = 2 0 (1) Find the radius of
Answer #1

ran Given equation is ayl + y + gry zo 6. multiply by :x , we get rid y t scy + gxy=0 , this is a Bessel equation of ordern=2 Ő = [ (antramy (8+n+) + «n (run) + 9 an 2) + 9,6132*1 = 7 An (+n] [v+n1) + (y (7+2)+gen.co ffor n=2,3 - 5 21(741) = 0 — =then n= am to (odd) then il so, we have, if n=2m (even an = am = (-9) m a. & (for m=1,2,.....) 22m (m) ( a = - 9 ao = (-9) am 22m (m12 m=1 onere cm - mio m= o so y = 2 +on+ go c. m => y= 2. [+ = , Tom, chowe Jok) is the Beassel function of the firs